OpenStudy (anonymous):

complete the square. x^2-4x=0

OpenStudy (zzr0ck3r):

(x-2)^2 = 4

OpenStudy (mathstudent55):

Take half of the coefficient of the x term and square it. The coefficient of the x term is -4. Half of -4 is -2. The square of -2 is 4. Now add 4 to both sides. x^2 - 4x + 4 = 0 + 4 (x - 2)^2 = 4

OpenStudy (zzr0ck3r):

^^^so hard to explain that on here...good job

OpenStudy (mathstudent55):

thanks